OpenStudy (anonymous):

-4(3 - 2x) + 2x = 2x - 8

OpenStudy (anonymous):

distribute the -4 first

OpenStudy (luigi0210):

First what you do is distribute out that -4 into the (3-2x) giving you -12+8x+2x=2x-8. Next combine like terms. so add x's and numbers together: 10x-12=2x-8 Now move variables to one side and numbers to the other: 8x=4 Now divide and solve: x=1/2